[Brahms-l] Shift report 20050211 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 11 2005 - 08:16:59 EST
1. RHIC has been good, with beam all night except 30 min. between dumps.
2. took data for 5 hours.

The following runs completed (MRS@45, FS@3)
     
          tr3      tr6
----------------------------  
MRS@A1050, FS@B1723
r13085    111k       275k

MRS@A2250, FS@B1723
r13086      1k         3k
r13087     86k       643k
r13088     82k       553k
r13089     32k       175k

MRS@B2250, FS@B1723
r13093     6k         54k
r13094    21k        179k

Minimum bias run: r13090 (82k)
TPC pedestal run: r13091
non-TPC peds run: r13092

r13110 (MRS@B2250, FS@A1723) continues into next shift.

3. got a "lost communication" error from tpc1. Nothing serious though. Hiro fixed it. The D3 has been acting very strange in the last part of the shiftm, amongst other things not reacting to ramping orders. twice the readout boards have been replaced without success. Reinforcements are called upon. the problem will hopefully be solved soon.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 11 08:17:22 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 11 2005 - 08:17:34 EST